# RRSIS: Referring Remote Sensing Image Segmentation

### v1

@@ Abstract
Localizing desired objects from remote sensing images is of great use in practical applications. Referring image segmentation, which aims at segmenting out the objects to which a given expression refers, has been extensively studied in natural images. However, almost no research attention is given to this task of remote sensing imagery. Considering its potential for real-world applications, in this paper, we introduce referring remote sensing image segmentation (RRSIS) to fill in this gap and make some insightful explorations. Specifically, we create a new dataset, called RefSegRS, for this task, enabling us to evaluate different methods. Afterward, we benchmark referring image segmentation methods of natural images on the RefSegRS dataset and find that these models show limited efficacy in detecting small and scattered objects. To alleviate this issue, we propose a language-guided cross-scale enhancement (LGCE) module that utilizes linguistic features to adaptively enhance multi-scale visual features by integrating both deep and shallow features. The proposed dataset, benchmarking results, and the designed LGCE module provide insights into the design of a better RRSIS model. We will make our dataset and code publicly available.
@@
In recent years, the volume of remote sensing data has grown considerably. Many tasks that rely solely on image processing and analysis, such as image classification, object detection, and semantic segmentation, have achieved impressive results [1]. However, models designed for these tasks only address a pre-determined set of categories, rather than understanding free-form expressions and specifically identifying regions of interest to end users. This may lead to a less targeted and suboptimal user experience. Thus, allowing end users to specify segmentation results for certain regions according to their individual needs would improve the efficiency and user interactivity of remote sensing image interpretation. In this work, we introduce referring remote sensing image segmentation (RRSIS) to provide an intuitive way for end users to specify objects of interest and enable targeted image analysis. Explicitly, given a remote sensing image and a language expression, RRSIS is to provide a pixel-level mask of desired objects based on the content of the image and the expression. Although referring image segmentation has been extensively studied in natural images, almost no attention has been given to this task in the context of remote sensing images. Thus, it is necessary to explore RRSIS, which enables users without expertise in the domain to obtain the precise information they require.
@@
Since RRSIS is a new task, and there are no relevant datasets available to provide segmentation masks for it, we create a new dataset called RefSegRS on top of images and pixel-wise annotations from the SkyScapes dataset [3]. RefSegRS dataset is constructed by designing various referring expressions and generating the corresponding segmentation masks automatically. The expressions include categories, attributes, or spatial relationships with other entities, as these are the features that end-users often use to refer to objects. The dataset consists of 4,420 image-language-label triplets. Fig. 1 shows two examples (natural images) from the VGPhraseCut dataset [2] and two examples (remote sensing images) from the proposed dataset. Generally speaking, in natural images, targets are prominent, and scenes are small-scale, whereas this is different in remote sensing images. Due to their top-down view and limited resolution, remote sensing imagery usually contains a wide range of object categories and relatively small objects.
@@
In order to comprehensively explore the task of RRSIS, we evaluate a range of referring image segmentation methods [4],[5], [6], [7], [8] on the RefSegRS dataset, which are originally designed for natural images. Our experiments indicate that directly applying existing referring image segmentation models to the RefSegRS dataset does not yield satisfactory results, as we find that small and scattered objects are often missed in the results. The reason behind this is that in natural images, objects tend to be large in size, while in remote sensing images, ground objects typically occupy fewer pixels and appear small and scattered, such as cars and road markings. Therefore, to successfully identify these hard samples, it is critical to design a dedicated deep network architecture for RRSIS tasks. Transformer networks can model long-range dependencies between different parts of an image and have an advantage in detecting scattered objects (such as scattered vehicles) and identifying correlations between different object parts (such as white stripes on a zebra crossing). Thus, to address the above-mentioned issue, we base our model on language-aware vision Transformer (LAVT) [8], which is currently considered one of the state-of-the-art for referring image segmentation tasks.
@@
Although LAVT can largely outperform existing convolutional neural network (CNN)-based methods, experimental results show that it still struggles with segmenting small and scattered objects in remote sensing images. For this issue, in this paper, we propose a Transformer-based feature enhancement module. It is motivated by two key points regarding characteristics of deep and shallow features. On the one hand, deep features exhibit high abstraction, and each pixel has strong expressive ability, which makes it easy to categorize them into different classes. On the other hand, while shallow features have high spatial resolution, making it possible to identify small objects, their semantic level is relatively low. Taking into account the above two points, we design a module to integrate deep and shallow features by leveraging linguistic features as guidance to better segment small objects.
@@ II-A Referring Image Segmentation on Natural Images
Typically, a referring image segmentation method first extracts visual and linguistic features from images and language expressions and then fuses the multimodal features to predict segmentation masks [8]. The pioneer work can be found in [4], where a long short-term memory (LSTM) network is employed to encode referring expressions into vectors, and a CNN is used to learn visual features. Then a fully convolutional network (FCN) generates segmentation masks based on both modalities. This inspires subsequent works, such as using DeepLab [9],[5] to learn better image representations and utilizing simple recurrent units (SRUs) [10] or bi-directional gated recurrent units (GRUs) [11] to extract better linguistic features from texts. In terms of the multimodal feature fusion, a range of attention mechanisms [12], [6], [7], [13] are studied to model the relationship between language and image modalities. Besides, inspired by the success of contrastive language-image pretraining (CLIP) [14], the authors of [15] propose a CLIP-driven framework for transferring knowledge obtained from CLIP to achieve text-to-pixel alignment. Transformers have recently showcased significant accomplishments in computer vision tasks. [8] introduces a Transformer network to achieve cross-modal alignment through the early fusion of language and visual features.
@@ II-C Visual Question Answering for Remote Sensing Data
Given an image and a corresponding question, a VQA system aims to provide a correct answer to the question based on the image. Lobry et al. [19] first introduce VQA into remote sensing by creating two datasets for this task. They use a CNN and a recurrent neural network (RNN) for feature extraction and then fuse multi-modal features to predict answers. Subsequently, a large-scale dataset named RSVQAxBEN [20] is established on top of the BigEarthNet dataset [21] to further advance the research on this task. To enhance the multi-modal representation, many methods have been investigated. The work [22] employs an attention mechanism and a bilinear technique to improve the alignment between visual and textual modalities. The paper [23] designs a bi-modal Transformer-based method to learn powerful multi-modal features. In [24], a spatial hierarchical reasoning network is designed toward a better fusion of visual and language features. Chappuis et al. [25] propose to extract context information from the visual input and inject it into a language model to obtain answers. Their method improves the performance as well as the interpretability of VQA systems. In addition, recent work [26] proposes an easy-to-hard training strategy to train VQA models. To improve the robustness of the VQA system, a multilingual augmentation method is designed in [27]. The work [28] introduces a change detection-based VQA model to extend this task to a multi-temporal domain.
@@ II-D Remote Sensing Image Captioning
Image captioning aims to describe the given images. [29] first introduces this task in remote sensing and utilizes a CNN and an RNN to extract visual features and generate words, respectively. Subsequently, [30] provides a comprehensive review of using handcrafted and deep features in this task. Encoder-decoder architectures are used in many works [31], [32], [33], where an encoder learns visual features from the input images and a decoder transfers the features to textual descriptions. In addition, various attention-based methods are studied a lot in recent years. For example, Zhao et al [34] develop a structured attention-based method, in order to use structural features in semantic information. Zhang et al. [35] propose a multi-source interactive attention network to investigate the impact of the semantic attribute and contextual information of generated words in producing descriptions. More recently, Transformers are also explored in this task. [36] exploits a Transformer network as an encoder-decoder and an auxiliary decoder to address the limited data issue. A multilayer aggregated Transformer is designed in [37] to leverage multi-scale features in the decoder.
@@ III-A Image Collection
We collect images by cropping large tiles from the SkyScapes dataset [3], which consists of 16 RGB tiles. Each tile has a size of pixels with a spatial resolution of 13 cm. All pixels are divided into 20 categories: low vegetation, paved road, non-paved road, paved parking place, non-paved parking place, bikeway, sidewalk, entrance/exit, danger area, lane marking, building, car, trailer, van, truck, large truck, bus, clutter, impervious surface, and tree. For those labeled as lane marking, the creators also offer their specific types such as dash-line, long-line, and small dash-line, but these labels are not considered in our case. To obtain information-rich images, we crop the tiles into images of pixels using a sliding window with a stride of 600 pixels. Considering the input size of deep neural networks, we further downsample them into pixels. Figure 2 shows several example images from the RefSegRS dataset.
@@ III-B Referring Expression Generation
Category or category with attribute. This is a commonly used expression for common users to indicate their interests in certain objects. Based on this template, we generate referring expressions by 1) directly specifying categories such as “vehicle” and “road” or 2) adding attributes to them, e.g., “light-weight vehicle” and “paved road”. After analyzing the SkyScapes dataset, we identify 14 distinct categories and 5 attribute tags that can be used to describe objects, as shown in Table I.
@@
The word cloud of the generated expressions is depicted in Fig. 3.
@@ III-C Mask Generation
In order to generate pixel-wise annotations for language expressions, we take the following steps:
@@
Step 2. We generate binary masks for all categories in the Skyscapes dataset and perform algebraic operations on them based on their conceptual relationships with language expressions. Masks corresponding to expressions generated by the first template can be generated at this step.
Step 3. For expressions generated using spatial relationships, we segment masks obtained from Step 2 into instance-level annotations and discard those not satisfying specified spatial relationships. We also add buffers to these instances in determining whether they are adjacent or not, so that outputs are more in line with human intuition visually. Afterward, the remaining instances are gathered for producing the final masks of expressions.
@@ III-D Dataset Statistics
The proposed dataset consists of 4,420 image-language-label triplets across 285 scenes. There are 151 scenes with 2,172 referring expressions in the training set, 31 scenes with 431 expressions in the validation set, and 103 scenes with 1,817 expressions in the test set. Note that referring expressions associated with the same scene are grouped into one set, which means that there is no overlap among the three subsets.
@@ IV Methodology
Referring image segmentation aims at segmenting out target objects with the guidance of natural language. To this end, representation learning, multi-modal feature fusion, and semantic segmentation are crucial components of this system. A typical referring image segmentation model usually consists of four parts: a visual encoder, a language encoder, a multi-modal feature fusion module, and a segmentation head. Specifically, in the stage of feature learning, visual features of a given image are extracted by the visual encoder, and linguistic features of the input text are computed by the language encoder. Subsequently, the multi-modal fusion module is used to integrate information from both vision and language modalities via a sophisticated fusion mechanism. In the end, the multi-modal features are fed into the segmentation head to predict the final segmentation map. By doing so, the corresponding objects described in the referring expression can be outlined by the model.
Thanks to the long-distance spatial modeling and multi-modal learning abilities of Transformers [38], [39], the performance of pixel-level dense prediction tasks has been greatly enhanced in both computer vision and remote sensing communities. Thus, we base our baseline method on a Transformer-based model LAVT [8]. Further, we design an LGCE module to improve the identification of small objects and scattered distributed objects in remote sensing images.
@@ IV-A Overall Architecture
We adopt LAVT as our baseline framework. Following it, we use Swin Transformer [40] as the backbone network for extracting visual features from the input image. For linguistic features, BERT [41] is utilized to encode the input language expression into feature vectors. Regarding multi-modal fusion, LAVT designs a pixel-word attention module (PWAM) and a language gate (LG) module to fuse language and visual features.
For the part of feature decoding, LAVT exploits a UNet-like architecture to enhance feature maps in the decoding stage with the encoded features by using simple residual connections. Although it has been proven to be effective for natural images, we argue that this simple combination of features with different spatial sizes is not optimal for remote sensing images. We observe that expressions accompanying such images often contain information about the size of target objects. For example, texts containing car and road marking tend to refer to small objects, while texts containing building generally relate to large objects. The accurate identification of small objects requires a more sophisticated multi-scale feature fusion module that takes the input text expressions into account. To this end, we design a language-guided Transformer module that enhances visual representations. As the proposed method is built upon a Transformer architecture, to clearly explain our approach, we first briefly describe Vision Transformer (ViT) [39] in the following subsection.

### v2

@@ Abstract
Localizing desired objects from remote sensing images is of great use in practical applications. Referring image segmentation, which aims at segmenting out the objects to which a given expression refers, has been extensively studied in natural images. However, almost no research attention is given to this task of remote sensing imagery. Considering its potential for real-world applications, in this paper, we introduce referring remote sensing image segmentation (RRSIS) to fill in this gap and make some insightful explorations. Specifically, we create a new dataset, called RefSegRS, for this task, enabling us to evaluate different methods. Afterward, we benchmark referring image segmentation methods of natural images on the RefSegRS dataset and find that these models show limited efficacy in detecting small and scattered objects. To alleviate this issue, we propose a language-guided cross-scale enhancement (LGCE) module that utilizes linguistic features to adaptively enhance multi-scale visual features by integrating both deep and shallow features. The proposed dataset, benchmarking results, and the designed LGCE module provide insights into the design of a better RRSIS model. The dataset and code will be available at https://gitlab.lrz.de/ai4eo/reasoning/rrsis.
@@
In recent years, the volume of remote sensing data has grown considerably. The surge in data volume has prompted the development of various techniques aiming at automatic perception of remote sensing imagery. During the early stages, traditional approaches, such as bag-of-words-based methods [1], are widely used in tasks including image classification [2] and object detection [3]. Additionally, unanimously accepted classical methods, including JSEG [4], graph cuts [5], watershed [6], superpixels [7], and EM-based algorithms [8], are developed and actively applied to the segmentation of remote sensing images.
With the advent of deep learning, numerous models based on deep neural networks have been proposed, achieving dominating performance on the perception of remote sensing images. Convolutional neural networks (CNNs) such as VGG [10], ResNet [11], and ConvNext [12] achieve human-competitive accuracy in image classification on large-scale datasets. More recently, Transformer-based networks like Vision Transformer (ViT) [13] and Swin Transformer [14], introduce the capacity to model long-distance relations within images, further enhancing the image classification performance. Moreover, object detection in remote sensing images has witnessed a dramatic improvement in accuracy with the emergence of models like Faster R-CNN [15], YOLO [16], and DETR [17]. Simultaneously, the task of semantic segmentation has also benefited immensely from deep learning techniques, with models like FCN [18], U-Net [19], Deeplab [20], and Segformer [21] attaining remarkable performance, allowing for pixel-level interpretation of remote sensing imagery.
@@
The models tailored for the tasks mentioned above categorize elements at the levels of images, objects, or pixels, without reliance on natural language guidance. This may lead to a less targeted and suboptimal user experience. Thus, allowing end users to specify segmentation results for certain regions according to their individual needs would improve the efficiency and user interactivity of remote sensing image interpretation. In this work, we introduce referring remote sensing image segmentation (RRSIS) to provide an intuitive way for end users to specify objects of interest and enable targeted image analysis. Explicitly, given a remote sensing image and a language expression, RRSIS is to provide a pixel-level mask of desired objects based on the content of the image and the expression. Although referring image segmentation has been extensively studied in natural images [22, 23, 24, 25, 26, 27, 28], almost no attention has been given to this task in the context of remote sensing images. Thus, it is necessary to explore RRSIS, which enables users without expertise in the domain to obtain the precise information they require.
@@
Derived from the concept of referring natural image segmentation, we introduce RRSIS as a novel task within the domain of remote sensing. Since there are no relevant datasets available for remote sensing data, we create a new dataset called RefSegRS on top of images and pixel-wise annotations from the SkyScapes dataset [29]. RefSegRS dataset is constructed by designing various referring expressions and generating the corresponding masks automatically. The expressions include categories, attributes, or spatial relationships with other entities, as these are the features that end-users often use to refer to objects. The dataset consists of 4,420 image-language-label triplets. Fig. 1 shows two examples (natural images) from the VGPhraseCut dataset [9] and two examples (remote sensing images) from the proposed dataset. Generally speaking, in natural images, targets are prominent, and scenes are small-scale, whereas this is different in remote sensing images. Due to their top-down view and limited resolution, remote sensing imagery usually contains a wide range of object categories and relatively small objects.
@@
In order to comprehensively explore the task of RRSIS, we evaluate a range of referring image segmentation methods [22, 23, 24, 25, 26] on the RefSegRS dataset, which are originally designed for natural images. Our experiments indicate that directly applying existing referring image segmentation models to the RefSegRS dataset does not yield satisfactory results, as we find that small and scattered objects are often missed in the results. The reason behind this is that in natural images, objects tend to be large in size, while in remote sensing images, ground objects typically occupy fewer pixels and appear small and scattered, such as cars and road markings. Therefore, to successfully identify these hard samples, it is critical to design a dedicated deep network architecture for RRSIS tasks. Transformer networks can model long-range dependencies between different parts of an image and have an advantage in detecting scattered objects (such as scattered vehicles) and identifying correlations between different object parts (such as white stripes on a zebra crossing). Thus, to address the above-mentioned issue, we base our model on language-aware vision Transformer (LAVT) [26], which is currently considered one of the state-of-the-art methods for referring image segmentation tasks.
@@
Although LAVT can largely outperform existing CNN-based methods, experimental results show that it still struggles with segmenting small and scattered objects in remote sensing images. For this issue, in this paper, we propose a Transformer-based feature enhancement module. It is motivated by two key points regarding characteristics of deep and shallow features. On the one hand, deep features exhibit high abstraction, and each pixel has strong expressive ability, which makes it easy to categorize them into different classes. On the other hand, while shallow features have high spatial resolution, making it possible to identify small objects, their semantic level is relatively low. Taking into account the above two points, we design a module to integrate deep and shallow features by leveraging linguistic features as guidance to better segment small objects.
@@ II-A Referring Image Segmentation on Natural Images
Typically, a referring image segmentation method first extracts visual and linguistic features from images and language expressions and then fuses the multimodal features to predict masks [30]. The pioneer work can be found in [22], where a long short-term memory (LSTM) [31] network is employed to encode referring expressions into vectors, and a CNN is used to learn visual features. Then a fully convolutional network (FCN) generates masks based on both modalities. This inspires subsequent works, such as using DeepLab [23], [30] to learn better image representations and utilizing simple recurrent units (SRUs) [32] or bi-directional gated recurrent units (GRUs) [33] to extract better linguistic features from texts. In terms of the multimodal feature fusion, a range of attention mechanisms [24], [25], [34], [35] are studied to model the relationship between language and image modalities. Besides, inspired by the success of contrastive language-image pretraining (CLIP) [36], the authors of [37] propose a CLIP-driven framework for transferring knowledge obtained from CLIP to achieve text-to-pixel alignment.
Transformers have recently showcased significant accomplishments in computer vision tasks. LAVT [26] employs bidirectional encoder representations from Transformers (BERT) [38] and a novel Transformer network for referring image segmentation. BERT [38], a model designed for extracting language representation, is pre-trained on a considerable volume of text data. It excels at learning features from text and can be fine-tuned for various downstream tasks, showing improvements in natural language processing-related tasks. Additionally, LAVT introduces a pixel-word attention module (PWAM), which aims at aligning visual representations with language features. Specifically, given visual and language features, PWAM performs convolutions, elements-wise multiplication, matrix multiplication, and softmax to align and fuse multi-modal features.
@@ II-C Visual Question Answering for Remote Sensing Data
Given an image and a corresponding question, a VQA system aims to provide a correct answer to the question based on the image. Lobry et al. [42] first introduce VQA into remote sensing by creating two datasets for this task. They use a CNN and a recurrent neural network (RNN) for feature extraction and then fuse multi-modal features to predict answers. Subsequently, a large-scale dataset named RSVQAxBEN [43] is established on top of the BigEarthNet dataset [44] to further advance the research on this task. To enhance the multi-modal representation, many methods have been investigated. The work [45] employs an attention mechanism and a bilinear technique to improve the alignment between visual and textual modalities. The paper [46] designs a bi-modal Transformer-based method to learn powerful multi-modal features. In [47], a spatial hierarchical reasoning network is designed toward a better fusion of visual and language features. Chappuis et al. [48] propose to extract context information from the visual input and inject it into a language model to obtain answers. Their method improves the performance as well as the interpretability of VQA systems. In addition, the work [49] introduces a change detection-based VQA model to extend this task to a multi-temporal domain.
@@ II-D Remote Sensing Image Captioning
Image captioning aims to describe the given images. [50] first introduces this task in remote sensing and utilizes a CNN and an RNN to extract visual features and generate words, respectively. Subsequently, [51] provides a comprehensive review of using handcrafted and deep features in this task. Encoder-decoder architectures are used in multiple works [52, 53, 54, 55], where an encoder learns visual features from the input images and a decoder transfers the features to textual descriptions. In addition, various attention-based methods are studied a lot in recent years. For example, Zhao et al. [53] develop a structured attention-based method, in order to use structural features in semantic information. Zhang et al. [56] propose a multi-source interactive attention network to investigate the impact of the semantic attribute and contextual information of generated words in producing descriptions. More recently, Transformers are also explored in this task. [57] exploits a Transformer network as an encoder-decoder and an auxiliary decoder to address the limited data issue. A multilayer aggregated Transformer is designed in [58] to leverage multi-scale features in the decoder.
@@ III-A Image Collection
We collect images by cropping large tiles from the SkyScapes dataset [29], which consists of 16 RGB tiles. Each tile has a size of pixels with a spatial resolution of 13 cm. All pixels are divided into 20 categories: low vegetation, paved road, non-paved road, paved parking place, non-paved parking place, bikeway, sidewalk, entrance/exit, danger area, lane marking, building, car, trailer, van, truck, large truck, bus, clutter, impervious surface, and tree. For those labeled as lane marking, the creators also offer their specific types such as dash-line, long-line, and small dash-line, but these labels are not considered in our case. To ensure that each image contains adequate objects and inter-object relations, we crop the tiles into images of pixels using a sliding window with a stride of 600 pixels. Considering the input size of deep neural networks, we further downsample them into pixels. Fig. 2 shows several example images from the RefSegRS dataset.
@@ III-B Referring Expression Generation
Category or category with attribute. This is a commonly used expression for common users to indicate their interests in certain objects. Based on this template, we generate referring expressions by 1) directly specifying categories such as “vehicle” and “road” or 2) adding attributes to them, e.g., “light-weight vehicle”. After analyzing the SkyScapes dataset, we identify 14 distinct categories and 5 attribute tags that can be used to describe objects, as shown in Table I.
@@
To offer a comprehensive visualization of the primary expressions within the dataset, we present a word cloud of the generated expressions in Fig. 3. As shown in it, “parking area” emerges as the most frequently recurring object in the dataset.
@@ III-C Mask Generation
Segmentation maps are sourced from the SkyScapes dataset [29], with each pixel in the maps indicating a class label. They are used to automatically generate ground truth masks (binary masks) based on natural language expressions for referring image segmentation tasks. In order to generate pixel-wise annotations for language expressions, we take the following steps:
@@
Step 2. To generate a binary mask for each category, we make use of pixel-wise annotations in the SkyScapes dataset and set values of pixels within/outside the category to 1/0. For categories composed of several subcategories (e.g., “vehicle” includes “car”, “van”, “truck”, etc.), we perform element-wise additions on binary masks of subcategories automatically.
Step 3. For expressions generated using spatial relationships, we partition masks obtained from step 2 into multiple instances (i.e., connected regions) and subsequently remove instances that do not conform to the specified spatial relationships. We also add buffers to these instances in determining whether they are adjacent or not, so that outputs are more in line with human intuition visually. Afterward, the remaining instances are gathered for producing the final masks of expressions.
@@ III-D Dataset Statistics
We conduct spatial relation and attribute analysis on semantic labels of the SkyScapes dataset to generate natural language expressions and the corresponding masks. After manually filtering out uninformative image-language-label triplets, we ultimately obtain a dataset consisting of 4,420 image-language-label triplets across 285 scenes. There are 151 scenes with 2,172 referring expressions in the training set, 31 scenes with 431 expressions in the validation set, and 103 scenes with 1,817 expressions in the test set. Note that referring expressions associated with the same scene are grouped into one set, which means that there is no overlap among the three subsets.
@@ IV Methodology
Referring image segmentation aims at segmenting out target objects with the guidance of natural language. To this end, representation learning, multi-modal feature fusion, and semantic segmentation are crucial components of this system. A typical referring image segmentation model usually consists of four parts: a visual encoder, a language encoder, a multi-modal feature fusion module, and a segmentation head [22]. Specifically, in the stage of feature learning, visual features of a given image are extracted by the visual encoder, and linguistic features of the input text are computed by the language encoder. Subsequently, the multi-modal fusion module is used to integrate information from both vision and language modalities via a sophisticated fusion mechanism. In the end, the multi-modal features are fed into the segmentation head to predict the final mask. By doing so, the corresponding objects described in the referring expression can be outlined by the model.
Thanks to the long-distance spatial modeling and multi-modal learning abilities of Transformers [59, 13], the performance of pixel-level dense prediction tasks has been greatly enhanced in both computer vision and remote sensing communities. Thus, we base our baseline method on a Transformer-based model LAVT [26]. Further, we design an LGCE module to improve the identification of small objects and scattered distributed objects in remote sensing images.
@@ IV-A Overall Architecture
We adopt LAVT as our baseline framework. Following it, we use Swin Transformer [14] as the backbone network for extracting visual features from the input image. Since language has a guiding function in achieving final results, it is necessary to learn semantic representations of language expressions. Hence, we incorporate BERT [38] to extract representative language embeddings. BERT is trained using two methods: 1) masked language modeling, predicting randomly masked words in a sentence; 2) next sentence prediction, determining if two sentences logically follow each other. This dual-training approach empowers BERT to understand complex languages, making it effective for diverse natural language processing tasks. In RRSIS tasks, BERT is employed for three advantages. First, BERT can help handle ambiguity and variability in natural language inputs. The representative language encoding enables the model to capture and understand the semantic meaning in the given language guidance. Second, the utilization of language embeddings from BERT offers scalability and flexibility. It allows the model to understand and process a wide range of language inputs and generalize to unseen input sentences. Third, the pre-trained language embeddings from BERT can enhance the efficiency of model training and improve performance. Regarding multi-modal fusion, LAVT designs a PWAM and a language gate (LG) module to fuse language and visual features.
For the part of feature decoding, LAVT exploits a UNet-like architecture to enhance feature maps in the decoding stage with the encoded features by using simple residual connections. Although it has been proven to be effective for natural images, we argue that this simple combination of features with different spatial sizes is not optimal for remote sensing images. We observe that expressions accompanying such images often contain information about the size of target objects. For example, texts containing car and road marking tend to refer to small objects, while texts containing building generally relate to large objects. The accurate identification of small objects requires a more sophisticated multi-scale feature fusion module that takes the input text expressions into account. To this end, we design a language-guided Transformer module that enhances visual representations. As the proposed method is built upon a Transformer architecture, to clearly explain our approach, we first briefly describe ViT [13] in the following subsection.